### v1

@@ 2 Mathematical model
where the velocity may take on both positive and negative values.
@@ 3 Numerical schemes
The numerical discretization to solve the equation (8) is done in space and time using the notation $t^{n}=n\tau$, $n=0,1,...N$ for a chosen $N$ and $\tau>0$, and the spatial discretization is based on the finite volume method [leveque_finite_2004]. The computational domain is divided into finite volumes with a regular grid in the form $I_{i}=(x_{i-1/2},x_{i+1/2})$, where $x_{i-1/2}=x_{L}+ih-h/2$ with $i,j=1,2,...,M$ for the chosen $M$ and $h=(x_{R}-x_{L})/M$.
To obtain the numerical scheme, we use the finite volume method, with the numerical approximation of space and time averaged values $\bar{q}_{i}$ and $\bar{u}_{i+1/2}^{n}$ as in, e.g., [shu_essentially_1998, leveque_finite_2004],
@@ 3.1 First order schemes
$+$ $v_{i+1/2}^{-}U_{i+1}^{n}-v_{i-1/2}^{-}U_{i}^{n}\Big{]}\,.$ (13)
@@
$F(U_{i}^{n+1})-F(U_{i}^{n})+\frac{\tau}{h}\Big{[}$ $v_{i+1/2}^{+}U_{i}^{n+1}-v_{i-1/2}^{+}U_{i-1}^{n+1}$
$+$ $v_{i+1/2}^{-}U_{i+1}^{n+1}-v_{i-1/2}^{-}U_{i}^{n+1}\Big{]}=0\,.$ (14)
@@
$i=1,\dots,M,$ (15)
$i=M,\dots,1.$ (16)
@@ 3.2 Second order accurate scheme
with the approximations $\partial_{x}U_{i}^{n+1}$ and $\partial_{t}U_{i}^{n+1}$ expressed as a linear combination of two different approximation choices [frolkovic2023high] using parameter $\boldsymbol{\omega}\in[0,1]$, which can be uniquely defined for each finite volume $I_{i}$ as $\omega_{i}$. Consequently, we obtain the compact implicit numerical approximation
$U_{i+1/2}^{n+1/2}=U_{i}^{n+1}$ $+\frac{1}{2}\big{(}\omega_{i}(U_{i}^{n+1}-U_{i-1}^{n+1})+(1-\omega_{i})(U_{i+1}^{n+1}-u_{i}^{n+1})\big{)}$
$-\frac{1}{2}\big{(}\omega_{i}(U_{i}^{n+1}-U_{i}^{n})+(1-\omega_{i})(U_{i+1}^{n+1}-U_{i+1}^{n})\big{)}$
$=U_{i}^{n+1}$ $-\frac{1}{2}\big{(}\omega_{i}(U_{i-1}^{n+1}-U_{i}^{n})+(1-\omega_{i})(U_{i}^{n+1}-U_{i+1}^{n})\big{)}\,.$ (19)
@@ 4 Two dimensional case
$u(x_{L},y,t)=u_{x_{L}}(y,t),\quad u(x,y_{L},t)=u_{y_{L}}(x,t),\quad u(x_{R},y,t)=u_{x_{R}}(y,t),\quad u(x,y_{R},t)=u_{y_{R}}(x,t)\,.$ (23)
For simplicity of the notation, we assume a squared computational domain that is divided into finite volumes of the form $I_{i,j}=(x_{i-1/2},x_{i+1/2})\times(y_{j-1/2},y_{j+1/2})$ with the regular square grid, where $x_{i-1/2}=x_{L}+ih-h/2$ and $y_{j-1/2}=x_{L}+jh-h/2$ with $i,j=1,2,...,M$ for the chosen $M$ and $h=(x_{R}-x_{L})/M$.
@@
$w=w^{+}+w^{-},\hskip 28.45274pt$ $\text{with }\,w^{+}\coloneqq\max(0,w),\,\,w^{-}\coloneqq\min(0,w).$ (24)
@@
The high resolution compact implicit numerical scheme will look like
$Q_{i,j}^{n+1}-Q_{i,j}^{n}+\frac{\tau}{h}\Big{[}$ $v^{+}_{i+1/2,j}\big{(}U_{i,j}^{n+1}-\frac{l_{i,j}^{x,+}}{2}\big{(}\omega_{i,j}^{x,+}(U_{i-1,j}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{x,+})(U_{i,j}^{n+1}-U_{i+1,j}^{n})\big{)}\big{)}$
$+$ $w^{+}_{i,j+1/2}\big{(}U_{i,j}^{n+1}-\frac{l_{i,j}^{y,+}}{2}\big{(}\omega_{i,j}^{y,+}(U_{i,j-1}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{y,+})(U_{i,j}^{n+1}-U_{i,j+1}^{n})\big{)}\big{)}$
-- $v^{+}_{i-1/2,j}\big{(}U_{i-1,j}^{n+1}-\frac{l_{i-1,j}^{x,+}}{2}\big{(}\omega_{i-1,j}^{x,+}(U_{i-2,j}^{n+1}-U_{i-1,j}^{n})+(1-\omega_{i-1,j}^{x,+})(U_{i-1,j}^{n+1}-U_{i,j}^{n})\big{)}\big{)}$
-- $w^{+}_{i,j-1/2}\big{(}U_{i,j-1}^{n+1}-\frac{l_{i,j-1}^{x,+}}{2}\big{(}\omega_{i,j-1}^{y,+}(U_{i,j-2}^{n+1}-U_{i,j-1}^{n})+(1-\omega_{i,j-1}^{y,+})(U_{i,j-1}^{n+1}-U_{i,j}^{n})\big{)}\big{)}$
$+$ $v^{-}_{i+1/2,j}\big{(}U_{i+1,j}^{n+1}-\frac{l_{i+1,j}^{x,-}}{2}\big{(}\omega_{i+1,j}^{x,-}(U_{i+2,j}^{n+1}-U_{i+1,j}^{n})+(1-\omega_{i+1,j}^{x,-})(U_{i+1,j}^{n+1}-U_{i,j}^{n})\big{)}\big{)}$
$+$ $w^{-}_{i,j+1/2}\big{(}U_{i,j+1}^{n+1}-\frac{l_{i,j+1}^{y,-}}{2}\big{(}\omega_{i,j+1}^{y,-}(U_{i,j+2}^{n+1}-U_{i,j+1}^{n})+(1-\omega_{i,j+1}^{y,-})(U_{i,j+1}^{n+1}-U_{i,j}^{n})\big{)}\big{)}$
-- $v^{-}_{i-1/2,j}\big{(}U_{i,j}^{n+1}-\frac{l_{i,j}^{x,-}}{2}\big{(}\omega_{i,j}^{x,-}(U_{i+1,j}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{x,-})(U_{i,j}^{n+1}-U_{i-1,j}^{n})\big{)}\big{)}$
-- $w^{-}_{i,j-1/2}\big{(}U_{i,j}^{n+1}-\frac{l_{i,j}^{y,-}}{2}\big{(}\omega_{i,j}^{y,-}(U_{i,j+1}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{y,-})(U_{i,j}^{n+1}-U_{i,j-1}^{n})\big{)}\big{)}\Big{]}\,=0$ (28)
with the parameter $\boldsymbol{l}=(\boldsymbol{l}^{x,\pm},\boldsymbol{l}^{y,\pm})$, and $\boldsymbol{\omega}=(\boldsymbol{\omega}^{x,\pm},\boldsymbol{\omega}^{y,\pm})\in[0,1]$, different for each $I_{i,j}$. For a detailed derivation of the scheme and definitions of parameters, we refer to [AACEE_2023_Zakova, zakova_numerical_2024].
@@ 5.1 One dimensional experiment
Table 3: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (34), using the maximum Courant number $C_{max}=4$, with different choices of $p=1/4,1/2,3/4$.
Figure 5: The initial condition, the exact solution, numerical solution - first and second order (high resolution with WENO approximations) of $q$ and $u$, for $p=1/4$ (top) $p=1/2$ (middle) $p=3/4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=4$.
@@
Table 4: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (34), using the maximum Courant number $C_{max}=4$, with different choices of $p=5/4,3/2,7/4$.
Figure 6: The initial condition, the exact solution, the numerical solution - first and second order (high resolution with WENO approximations) of $q$ and $u$, for $p=5/4$ (top) $p=3/2$ (middle) $p=7/4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=4$.
@@
Table 5: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (34), using the maximum Courant number $C_{max}=4$, with different choices of $p=2,3,4$.
Figure 7: The initial condition, the exact solution, the numerical solution - first and second order (high resolution with WENO approximations) of $q$ and $u$, for $p=2$ (top) $p=3$ (middle) $p=4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=4$.
@@ 5.2 One dimensional experiment with variable velocity
The next experiment in one dimension demonstrates a visual comparison (Figures 8, 9) of the results using the variable velocity $v(x)=\cos(x)$, which can take on both positive and negative values, as described in equation (8); therefore, it is necessary to perform at least two sweeps (15)-(16) of the fast sweeping method. In this case, we choose $x\in[-4,11]$ and $T=1.5$ with $N=M/80$ ($M$ defined later) and $C_{max}=7.5$ (32).
@@
Figure 8: The comparison of the numerical solution ($q$ and $u$) first order (green color), high resolution second order with WENO approximation (red color), for $p=1/4$ for different choices of $M=320$ (top) $M=640$ (middle) $M=1280$ (bottom) for the final time $T=3$. The initial condition is shown as a gray dashed line.
@@
Figure 9: The comparison of the numerical solution ($q$ and $u$) first order (green color), high resolution second order with WENO approximation (red color), for $p=4$ for different choices of $M=320$ (top) $M=640$ (middle) $M=1280$ (bottom) for the final time $T=3$. The initial condition is shown as a gray dashed line.
@@ 6 Conclusion
Overall, this work not only contributes to the theoretical understanding of nonlinear transport phenomena but also provides practical numerical tools that can be utilized in various fields such as environmental hydrology and materials science. To offer a significant practical value of the numerical schemes presented, future research must explore their development for a system of coupled transport equations [donat_weno_2024, schmidt-traubPreparativeChromatography2020, frolkovicNumericalSimulationContaminant2016].

### v2

@@ 2 Mathematical model
where the velocity may take on both positive and negative values and the Dirichlet boundary condition will be extended also to the right boundary
$u(x_{R},t)=u_{x_{R}}(t)\,.$ (9)
@@ 3 Numerical schemes
The numerical discretization to solve the equation (8) is done in space and time using the notation $t^{n}=n\tau$, $n=0,1,...N$ for a chosen $N$ and $\tau>0$, and the spatial discretization is based on the finite volume method [leveque_finite_2004]. The computational domain is divided into finite volumes with a regular grid in the form $I_{i}=(x_{i-1/2},x_{i+1/2})$, where $x_{i-1/2}=x_{L}+(i-1)h$ with $i=1,2,...,M$ for the chosen $M$ and $h=(x_{R}-x_{L})/M$.
To obtain the numerical scheme, we use the finite volume method, with the numerical approximation of space and time averaged values $\bar{q}_{i}^{n}$ and $\bar{u}_{i+1/2}^{n+1/2}$ as in, e.g., [shu_essentially_1998, leveque_finite_2004],
@@ 3.1 First order schemes
$+$ $v_{i+1/2}^{-}U_{i+1}^{n}-v_{i-1/2}^{-}U_{i}^{n}\Big{]}\,,$ (14)
where the explicit structure of the scheme is clearly evident.
@@
$Q_{i}^{n+1}-Q_{i}^{n}+\frac{\tau}{h}\Big{[}$ $v_{i+1/2}^{+}U_{i}^{n+1}-v_{i-1/2}^{+}U_{i-1}^{n+1}$
$+$ $v_{i+1/2}^{-}U_{i+1}^{n+1}-v_{i-1/2}^{-}U_{i}^{n+1}\Big{]}=0\,,$ (15)
where the nonlinearity is readily apparent in the previously mentioned relation $Q_{i}^{n+1}=F(U_{i}^{n+1})$.
@@
First sweep: $i=1,\dots,M-1,$ (16)
Second sweep: $i=M-1,\dots,1.$ (17)
@@ 3.2 Second order accurate scheme
with the approximations $\partial_{x}U_{i}^{n+1}$ and $\partial_{t}U_{i}^{n+1}$ expressed as a linear combination of two different approximation choices [frolkovic2023high] using parameter $\boldsymbol{\omega}\in[0,1]$
$\partial_{x}U_{i}^{n+1}$ $=\frac{1}{h}\big{(}\omega(U_{i}^{n+1}-U_{i-1}^{n+1})+(1-\omega)(U_{i+1}^{n+1}-U_{i}^{n+1})\big{)}\,,$
$\partial_{t}U_{i}^{n+1}$ $=\frac{1}{\tau}\big{(}\omega(U_{i}^{n+1}-U_{i}^{n})+(1-\omega)(U_{i+1}^{n+1}-U_{i+1}^{n})\big{)}\,.$ (20)
Consequently, combining (19) with (20), we obtain the compact implicit numerical approximation
$U_{i+1/2}^{n+1/2}=U_{i}^{n+1}-\frac{1}{2}\big{(}\omega(U_{i-1}^{n+1}-U_{i}^{n})+(1-\omega)(U_{i}^{n+1}-U_{i+1}^{n})\big{)}\,.$ (21)
@@ 4 Two dimensional case
$u(x_{L},y,t)=u_{x_{L}}(y,t)\,,$ $\quad u(x,y_{L},t)=u_{y_{L}}(x,t)\,,$
$u(x_{R},y,t)=u_{x_{R}}(y,t)\,,$ $\quad u(x,y_{R},t)=u_{y_{R}}(x,t)\,.$ (25)
For simplicity of the notation, we assume a squared computational domain ($x,y\in(x_{L},x_{R})$) that is divided into finite volumes of the form $I_{i,j}=(x_{i-1/2},x_{i+1/2})\times(y_{j-1/2},y_{j+1/2})$ with the regular square grid, where $x_{i-1/2}=x_{L}+(i-1)h$ and $y_{j-1/2}=x_{L}+(j-1)h$ with $i,j=1,2,...,M$ for the chosen $M$ and $h=(x_{R}-x_{L})/M$.
@@
$w=w^{+}+w^{-},\hskip 28.45274pt$ $\text{with }\,w^{+}\coloneqq\max(0,w),\,\,w^{-}\coloneqq\min(0,w){\color[rgb]{0,0,0}\,,}$ (26)
@@
The high resolution compact implicit numerical scheme will then use the approximations for the values ${U}_{i+1/2,j}^{n+1/2}$ and ${U}_{i,j+1/2}^{n+1/2}$ in (4) as
${U}_{i+1/2,j}^{n+1/2}$ $=U_{i,j}^{n+1}-\frac{l_{i,j}^{x,+}}{2}\big{(}\omega_{i,j}^{x,+}(U_{i-1,j}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{x,+})(U_{i,j}^{n+1}-U_{i+1,j}^{n})\big{)}\,,$
${U}_{i,j+1/2}^{n+1/2}$ $=U_{i,j}^{n+1}-\frac{l_{i,j}^{y,+}}{2}\big{(}\omega_{i,j}^{y,+}(U_{i,j-1}^{n+1}-U_{i,j}^{n})+(1-\omega_{i,j}^{y,+})(U_{i,j}^{n+1}-U_{i,j+1}^{n})\big{)}\,,$ (30)
which stand next to the positive velocities $v^{+}_{i+1/2,j}$ and $w^{+}_{i,j+1/2}$; and the approximations ${U}_{i+1/2,j}^{n+1/2}$ and ${U}_{i,j+1/2}^{n+1/2}$ for negative cases
${U}_{i+1/2,j}^{n+1/2}$ $=U_{i+1,j}^{n+1}-\frac{l_{i+1,j}^{x,-}}{2}\big{(}\omega_{i+1,j}^{x,-}(U_{i+2,j}^{n+1}-U_{i+1,j}^{n})+(1-\omega_{i+1,j}^{x,-})(U_{i+1,j}^{n+1}-U_{i,j}^{n})\big{)}\,,$
${U}_{i,j+1/2}^{n+1/2}$ $=U_{i,j+1}^{n+1}-\frac{l_{i,j+1}^{y,-}}{2}\big{(}\omega_{i,j+1}^{y,-}(U_{i,j+2}^{n+1}-U_{i,j+1}^{n})+(1-\omega_{i,j+1}^{y,-})(U_{i,j+1}^{n+1}-U_{i,j}^{n})\big{)}\,,$ (31)
which stand next to the velocities $v^{-}_{i+1/2,j}$ and $w^{-}_{i,j+1/2}$, with the parameter $\boldsymbol{l}=(\boldsymbol{l}^{x,\pm},\boldsymbol{l}^{y,\pm})\in[0,1]$, and $\boldsymbol{\omega}=(\boldsymbol{\omega}^{x,\pm},\boldsymbol{\omega}^{y,\pm})\in[0,1]$, different for each $I_{i,j}$. For a detailed derivation of the scheme and definitions of parameters, we refer to [AACEE_2023_Zakova, zakova_numerical_2024].
@@ 5.1 One dimensional experiment
This can also be observed in Figures 4 for $C_{max}=0.25$ and in Figure 5 for $C_{max}=0.5$, which present the comparison of CPU time versus errors on a logarithmic scale.
Figure 4: CPU time versus errors in log scale for the first and second order accurate numerical scheme (explicit and implicit) with $C_{max}=0.25$.
Figure 5: CPU time versus errors in log scale for the first and second order accurate numerical scheme (explicit and implicit) with $C_{max}=0.5$.
@@
Table 3: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (37), using the maximum Courant number $C_{max}=6$, with different choices of $p=1/4,1/2,3/4$.
Figure 7: The initial condition, the exact solution, numerical solution - first and second order (high resolution with WENO approximations) - $q$ (left) and $u$ (right), for $p=1/4$ (top) $p=1/2$ (middle) $p=3/4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=6$.
@@
Table 4: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (37), using the maximum Courant number $C_{max}=6$, with different choices of $p=5/4,3/2,7/4$.
Figure 8: The initial condition, the exact solution, the numerical solution - first and second order (high resolution with WENO approximations) - $q$ (left) and $u$ (right), for $p=5/4$ (top) $p=3/2$ (middle) $p=7/4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=6$.
@@
Table 5: The numerical errors, and the EOCs using the first and second order (high resolution with WENO approximations) accurate implicit schemes, for a case with a discontinuous initial condition (37), using the maximum Courant number $C_{max}=6$, with different choices of $p=2,3,4$.
Figure 9: The initial condition, the exact solution, the numerical solution - first and second order (high resolution with WENO approximations) - $q$ (left) and $u$ (right), for $p=2$ (top) $p=3$ (middle) $p=4$ (bottom) with the final time $T=3$, $M=320$, $C_{max}=6$.
@@ 5.2 One dimensional experiment with variable velocity
The next experiment in one dimension demonstrates a visual comparison (Figures 10, 11) of the results using the variable velocity $v(x)=\cos(x)$, which can take on both positive and negative values, as described in equation (8); therefore, it is necessary to perform at least two sweeps (16)-(17) of the fast sweeping method. In this case, we choose $x\in[-4,11]$ and $T=1.5$ with $N=M/80$ ($M$ defined later) and $C_{max}=7.5$ (35).
@@
Figure 10: The comparison of the numerical solution ($q$ on the left side and $u$ on the right) first order (green color), high resolution second order with WENO approximation (red color), for $p=1/4$ for different choices of $M=320$ (top) $M=640$ (middle) $M=1280$ (bottom) for the final time $T=1.5$. The initial condition is shown as a gray dashed line.
@@
Figure 11: The comparison of the numerical solution ($q$ on the left side and $u$ on the right) first order (green color), high resolution second order with WENO approximation (red color), for $p=4$ for different choices of $M=320$ (top) $M=640$ (middle) $M=1280$ (bottom) for the final time $T=1.5$. The initial condition is shown as a gray dashed line.
@@ 6 Conclusion
Overall, this work contributes to the theoretical understanding of nonlinear transport phenomena by developing and analyzing compact implicit high resolution numerical schemes that effectively address the nonlinearities present in transport equations with sorption isotherms. In addition, it provides practical numerical tools applicable to various fields, such as environmental hydrology and materials science. The paper demonstrates that these schemes achieve higher accuracy without the restrictive time step limitations of explicit methods and that the use of limiters in high resolution schemes ensures physically realistic solutions even in the presence of discontinuities. To offer a significant practical value of the numerical schemes presented, future research must explore their development for a system of coupled transport equations [donat_weno_2024, schmidt-traubPreparativeChromatography2020, frolkovicNumericalSimulationContaminant2016].